Consider the following spot interest rates for maturities of one, two, three, and four years.

   r1 = 4.4%    r2 = 4.9%     r3 = 5.6%     r4 = 6.4%

Assuming a constant real interest rate of 2 percent, what are the approximate expected inflation rates for the next four years? (Do

l1 ___ %

l2 ___

l3 ___

l4____
Answer #1

Inflation rate in year 1 = 0.044 - 0.02 = 2.20%

Inflation rate in year 2 = 0.049 - 0.02 = 2.90%

Inflation rate in year 3 = 0.056 - 0.02 = 3.60%

Inflation rate in year 4= 0.064 - 0.02 = 4.40%